OpenStudy (anonymous):

Use substitution to solve the equation. 4x - y = 15 -2x + 3y = 12

OpenStudy (anonymous):

y=4x-15 -2x+3(4x-15)=12 -2x+12x-45=12 10x=57 x=57/10 y=39/5

Directrix (directrix):

4x - y = 15 Solve for y: -y = 15 - 4x y = -15 +4x -2x + 3y = 12 -2x + 3(-15 +4x) = 12 -2x -45 +12x = 12 10x = 57 x = 57/10 ** y = -15 +4x y = -15 + 4(57/10) = -15 + 114/5 = -75/5 + 114/5 = 39/5 *** (57/10 , 39/5) Decimals: (5.7, 7.8) *****